# Secure Information Embedding and Extraction in Forensic 3D Fingerprinting

### v1

@@ 4 SIDE Overview
The resemblance between G-code and assembly code motivates the use of TEE to protect the integrity of the printing process. Indeed, both G-code and assembly code are instructions for lower-level hardware, but both are oblivious to the inner workings of that hardware. For example, an instruction to fetch memory data to a register does not specify how this operation is implemented in the integrated circuit (IC). Meanwhile, two ICs that support the same instruction set may have completely different circuit designs.
For secure extraction of information in case of breaks, the information $\mathbf{w}$ goes through an encoding process before being embedded. The resulting codeword $\mathbf{c}$ has the $t$-break-resilient property; as the name suggests, the corresponding decoder is able to recover $\mathbf{w}$ even if the codeword is sliced into at most $(t+1)$ unordered fragments.
Next, the SIDE slicer generates G-code with an embedded break-resilient codeword using some method (e.g., varying layer height), and the SIDE firmware executes the G-code, i.e., translates it to signals that directly operate the printer’s actuators. For secure embedding purposes, both the generation and the execution of the G-code are isolated in a trusted environment. Then, the result is forwarded to the peripherals that communicate with printer actuators; this process is also assumed to be protected by TEE.
@@ 5 Secure Information Extraction
5.1 Overview
The secure information extraction feature of SIDE is attributed to a break-resilient code (BRC), which is specifically developed for forensic fingerprinting purposes. BRC includes an encoder which holds a binary string $\mathbf{w}\in{\{0,1\}}^{k}$ (an information word, e.g., printer ID) for some integer $k$, that is to be encoded to a string $\mathbf{c}\in{\{0,1\}}^{n}$ (a codeword). After encoding $\mathbf{w}$, the codeword $\mathbf{c}$ is embedded in a printed object. BRC may be applied to any possible bit embedding method, and yet in Section 6.2 we focus on varying layer width. For a security parameter $t$, an adversary breaks these $n$ bits at arbitrary $t$ locations or less, resulting in a multiset of at most $t+1$ fragments.
@@ 5.1.1 Systematic Reed-Solomon Codes
Systematic Reed-Solomon (RS) codes are a special type of Reed-Solomon codes, which have been widely employed in communication systems and data storage applications; for an introduction to the topic see [27, Ch. 5].
@@ 5.1.2 Run-length Limited Codes
A run-length limited (RLL) code is a collection of codewords in which the length of runs of repeated bits is bounded. Specifically, in this paper we employ the RLL code from [28, Algorithm 1] for its simplicity of implementation.
@@ 5.1.3 Mutually Uncorrelated Codes
A mutually uncorrelated (MU) code has the property that for every two (possibly identical) codewords, the prefix of one is not identical to the suffix of another. As such, the codewords of a MU code do not overlap with each other when appearing as substrings of a binary string. MU codes has been extensively investigated in the past [29, 30, 31, 32, 33, 34, 35, 36, 37, 28].
We adopt a classic construction of MU code, in which each codeword starts with ${\lceil\log k\rceil}$ zeros followed by a one, where $k$ is the length of the (binary) information word. The last bit is fixed to one, and the remaining bits are free from zero runs of length ${\lceil\log k\rceil}+1$.
5.1.4 Distinct Strings
Inspired by ideas from [28, Algorithm 1] and [38, Algorithm 1], we provide Algorithm 3 that maps an input word to an array of pairwise distinct binary strings, and offer the inverse operation in Algorithm 4. These procedures serve as an important component in BRC. Both algorithms, as well as proofs of their correctness, are given in the appendix.
Algorithm 1 Encode
@@ 5.1.4 Distinct Strings
5.2 Encoding
The encoding procedure of BRC is provided in Algorithm 1. Let $\mathbf{w}\in{\{0,1\}}^{k}$ be an information word, where
@@ 5.2 Encoding
and where $l$ and $m$ are positive integers such that $m\geq{\lceil 2\log l\rceil}+2$. The encoding of $\mathbf{w}$ begins by prepending it with $\mathbf{p}^{(0)},\mathbf{p}^{(1)},\cdots,\mathbf{p}^{(t-1)}$, resulting in
@@
where $\mathbf{p}^{(i)}$ is the binary representation of $i\in[0,t-1]$ in $m$ bits (line 3). Then, as shown in line 4, the resulting string $\mathbf{u}$ is fed into the function d-encode (which is implemented in Algorithm 3 in the appendix) and mapped to an array of $l$ pairwise distinct binary strings of length $m$, i.e.,
@@
$\texttt{next}[\mathbf{s}]=\begin{cases}\mathbf{u}_{i+1}\leavevmode\nobreak\ &\text{if}\leavevmode\nobreak\ \mathbf{s}=\mathbf{u}_{i}\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ i\in[0,l-2],\\ \mathbf{s}\leavevmode\nobreak\ &\text{otherwise.}\end{cases}$
Note that $\texttt{next}(\mathbf{s})$ is well defined, since $\mathbf{u}_{i}\neq\mathbf{u}_{j}$ for every $i\neq j$ by the pairwise distinct property of dStrings. It is also worth noting that the mapping from dStrings to next is injective; one may recover dStrings from next by observing every key $\mathbf{r}$ such that $\texttt{next}(\mathbf{r})\neq\mathbf{r}$, and connecting every two $\mathbf{r}_{a},\mathbf{r}_{b}$ of them if $\texttt{next}(\mathbf{r}_{a})=\mathbf{r}_{b}$.
@@
[…]
@@ 5.3 Decoding
The decoding algorithm may fail to identify a non-marker (i.e., $\mathbf{u}_{i}$ for $i\geq t$) in the information region if a break occurs inside it. Hence, the algorithm fails to identify at most $t_{1}$ non-markers. Notice that, failing to capture a non-marker $\mathbf{u}_{a}$ affects at most two key-value pairs of next, i.e., $\texttt{next}[\mathbf{u}_{a-1}]$ and $\texttt{next}[\mathbf{u}_{a}]$, and hence next and approxNext differ in at most $2t_{1}$ positions. ∎
The decoding algorithm proceeds to correct the constructed approxNext to next, i.e., the key-value store generated in Algorithm 1 from $\mathbf{w}$, using the collected redundancy strings and a standard Reed-Solomon decoder (line 10). Note that the missing redundancy strings are erasures (line 3), since their locations are known. The success of the decoding process is guaranteed in the following theorem.
@@
Line 10 outputs the correct key-value store next.
@@
Meanwhile, Lemma 3 implies that the number of key-value pairs in which next and approxNext differ is bounded by $2t_{1}$. Hence, the systematic part of the constructed codeword has at most $2t_{1}$ errors. Recall that the encoding process generated $4t$ redundancy strings (symbols in $\mathbb{F}_{2}^{m+1}$), and hence the decoding in line 10 is guaranteed to be successful since
@@
Finally, the array dStrings is obtained from next (line 11), and the function d-decode is employed to recover $\mathbf{u}$. Recall that $\mathbf{u}=\mathbf{p}^{(0)}\circ\cdots\circ\mathbf{p}^{(t-1)}\circ\mathbf{w}$, and hence the decoding procedure ends concludes by returning the $l\cdot m$ rightmost bits of $\mathbf{u}$ (line 13).
@@ 6 Implementation
We implemented a prototype of SIDE and conducted experiments to validate its applicability. First, the information bits $\mathbf{w}$ are encoded to a BRC codeword as described in Section 5.2. Then, we employ a Creality Ender 3 3D printer (Figure. 5(a)) to print objects with an embedded codeword (Figure 3(a)). The printer is operated by the Kipper firmware under the protection of OP-TEE. To read the information, a Leica S9D microscope is utilized to extract bits from the fragments of the printed object (Figure 3(b)). Note that reading externally using a visual device is not strictly necessary, as explained in the sequel. Finally, the bit strings that were recovered from the fragments go through the decoding process described in Section 5.3, and the original information is recovered.
6.1 BRC Encoding and Decoding
We implemented the BRC encoding and decoding procedures, whose pseudo-code is given in Algorithm 1 and Algoritm 2, respectively. The encoding part is written in C++ for the ease of using TEE. The implementation details and the use of TEE will be described in the future version of this paper. The decoding part is implemented in Python.
Figure 4: (a) A cuboid with an embedded BRC codeword. The codeword creates grooves on the outside of the object that are visible to the naked eye. For simplicity, we read the codeword using a microscope which detects light reflected from the top of each layer (Fig. 7). In cases where the outside grooves are not visible (e.g., scraping, filing, or coating), one can use a CT scanner to detect layer widths. (b) A cuboid broken by hand to $3$ fragments. Indeed, breaks occur naturally along cross-sections of the printing axis.
6.2 Information Embedding and Extraction
We argue that any practical bit embedding method should satisfy the following three properties.
@@ 6.2 Information Embedding and Extraction
The readability requirement implies economical reading of bits from the broken object with respect to a given use case. The robustness requirement states that the embedded information must be sufficiently robust against adversarial tempering, within a certain bound. For example, the surface of the printed object may be easily abraded and as a result, embedding information on the surface is not robust.
@@
The obliviousness requirements decouples the embedding method both from the information to be embedded and from the 3D models. Examples which violate the obliviousness requirement include representing $1$’s by thick layers and representing 00’s by thin layers; this will cause codewords with different number of 1’s to have embeddings of different lengths. Moreover, embedding bits by alerting the geometry of the printed object is also considered inappropriate, as it may significantly affect its functionality. Note that for any embedding method, imposing some requirement on the dimension of the 3D model is inevitable (e.g., that its size is sufficient for embedding a certain number of bits), and we do not consider this as a violation of the obliviousness requirement.
ptptptptptpt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
pt
$\mathrm{0.36\ mm}$
ptptpt
@@
Figure 5: Demonstration of embedding $01101001$ by varying layer width, in which each 00 is represented by three layers of $0.12$ millimeters, and each $1$ is represented by two layers of $0.18$ millimeters. The length required for embedding one bit is irrelevant to its value.
In our implementation we embed bits by varying layer width of the printed object, with a slight modification. As shown in Figure 5, a 00 is represented by three layers of $0.12$ millimeters, and a $1$ is represented by two layers of $0.18$ millimeters. These values are chosen to maximize information density, i.e., the number of bits that can be embedded per unit length, with respect to the minimum layer height ($0.12$ millimeters) offered by the Ender 3 printer used in the experiments. As the heights add up to $0.36$ millimeters in both cases, this method avoids the situation of height discrepancy mentioned in the previous paragraph, and satisfies the obliviousness requirement. As for the robustness requirement, we note that layer height is an intrinsic property of the printed object, and cannot be easily altered, nor it is affected by breaking or scraping the object.
The reading of information from fragments of the printed object is done by a Leica S9D microscope, which provides a high resolution that captures their rugged surface. It is worth noting that the reading of bits does not require the surface of the fragment to be free of adversarial tempering, as less-economical solutions, e.g., using an industrial CT scanner, can be used to infer the layer width via tomographic analysis.

### v3

@@ 4.1 Preliminary and overview
Furthermore, both G-code and assembly code are commands for lower-level hardware, but both are oblivious to the inner workings of that hardware. For example, a CPU instruction to fetch memory data to a register does not specify how this operation is implemented in the integrated circuit (IC). Meanwhile, two ICs that support the same instruction set may have completely different circuit designs. Examples of jeopardizing the compilation process have been shown in [thompson1984reflections]; with a deliberately crafted compiler, benign source code may be miscompiled to a malware. A similar attack has also been demonstrated in the recent backdoor implantation in the XZ Utils [xz].
4.2 Embedding Framework
Figure 7: Illustration of the interactions among the rich execution environment (REE), the trusted execution environment (TEE), and the physical world during the printing process. At first, the 3D model is fed into the host application (in REE), which invokes layer-gen (in TEE) to perform slicing and information embedding. For every bit of the to-be-embedded codeword $\mathbf{c}$, a G-code file is created by toolpath-gen and picked up by firmware (both in TEE). The firmware translates the G-code commands to stepper events, and invokes the hardware (stepper motors) to perform the actual printing in the physical world.
The SIDE embedding framework employs a host application that resides in the REE. It solely provides a frontend interface to the printer users, accepting 3D models and passing them to the backend, which wholly resides in TEE. The SIDE backend involves four major functional modules, namely codeclayer-gentoolpath-gen, and firmware.
At first, the codec encodes the information $\mathbf{w}$ to a break-resilient codeword $c$; the details of the encoding process will be given in Section 5. Next, the SIDE backend proceeds to the actual printing process. Recall that our embedding method relies on varying the thickness of layers. Specifically, it first segments the model into constant-length height intervals along the printing direction, and then associates the $i$-th height interval with the the $i$-th bit in $\mathbf{c}$. The value of $\mathbf{c}[i]$ determines the thickness of layers within the corresponding height interval (see Section 6.1 for details).
After picking up the model file from the host application, the following procedure is executed by the SIDE backend for every bit of $\mathbf{c}$ for secure information embedding. For every $\mathbf{c}[i]$layer-gen generates the cross-sectional diagrams of its associating layers. Next, layer-gen passes the diagrams to toolpath-gen along with their corresponding heights, i.e., their distances to the printer bed.
The toolpath-gen module then determines the nozzle toolpath from the received data, which is usually represented as a G-code file. Indeed, the G-code document for $\mathbf{c}[i]$ instructs the 3D printer on how to build the object within a certain height interval. Finally, the firmware module picks up the G-code document, and translates it to a series of stepper events. Through the board peripherals (GPIO), the SIDE backend passes these events to printer steppers, and conduct the actual printing in the physical world. This concludes the handling of $\mathbf{c}[i]$, and a graphical illustration is shown in Figure 7.
Note that the above procedure is repeated for every bit of the BRC codeword $\mathbf{c}$, and the printing process is concluded after handling all bits. Since the entire printing process is hidden in the trusted world and no intermediate data (e.g., a G-code file) is exposed to the user, the adversary is unable to strip off the embedded bits.
@@ 5 Secure Information Extraction
5.1 Preliminary and overview
The secure information extraction feature of SIDE is attributed to a break-resilient code (BRC), which is specifically developed for forensic fingerprinting purposes. Before being embedded, the to-be-embedded information $\mathbf{w}$ goes through an encoding process. The resulting codeword $\mathbf{c}$ has the $t$-break-resilient property; as the name suggests, the corresponding decoder is able to recover $\mathbf{w}$ even if the codeword is broken into at most $(t+1)$ unordered fragments.
BRC includes an encoder which holds a binary string $\mathbf{w}\in{\{0,1\}}^{k}$ (an information word, e.g., printer ID) for some integer $k$, that is to be encoded to a string $\mathbf{c}\in{\{0,1\}}^{n}$ (a codeword). After encoding $\mathbf{w}$, the codeword $\mathbf{c}$ is embedded in a printed object. BRC may be applied to any possible bit embedding method, and yet in Section 6.1 we focus on varying layer thickness. For a security parameter $t$, an adversary breaks these $n$ bits at arbitrary $t$ locations or less, resulting in a multiset of at most $t+1$ fragments [^3] For the flow of the paper, the proofs of theorems and lemmas are given in the appendix..
@@ 5.1.1 Systematic Reed-Solomon Codes
Systematic Reed-Solomon (RS) codes are a special type of Reed-Solomon codes, which have been widely employed in communication systems and data storage applications; for an introduction to the topic see [roth2006coding, Ch. 5].
@@ 5.1.2 Run-length Limited Codes
A run-length limited (RLL) code is a collection of codewords in which the length of runs of repeated bits is bounded. Specifically, in this paper we employ the RLL code from [levy2018mutually, Algorithm 1] for its simplicity of implementation.
@@ 5.1.3 Mutually Uncorrelated Codes
A mutually uncorrelated (MU) code has the property that for every two (possibly identical) codewords, the prefix of one is not identical to the suffix of another. As such, the codewords of a MU code do not overlap with each other when appearing as substrings of a binary string. MU codes has been extensively investigated in the past [Lev64, Lev70, Gilbert60, Bajic2004, Bajic2014, Chee2013, Bilotta2012, Blackburn2015, Wang2022, levy2018mutually]. In this paper, we adopt a classic construction of MU code, in which each codeword starts with ${\lceil\log k\rceil}$ zeros followed by a one, where $k$ is the length of the (binary) information word. The last bit is fixed to one, and the remaining bits are free from zero runs of length ${\lceil\log k\rceil}+1$.
5.1.4 Distinct Block Codes
Inspired by ideas from [levy2018mutually, Algorithm 1] and [repeat-free, Algorithm 1], we provide Alg. 3 that maps an input word to an array of distinct binary strings, and offer the inverse operation in Alg. 4. These procedures serve as an important component in BRC. Both algorithms, as well as proofs of their correctness, are given in the appendix.
5.2 Encoding
Algorithm 1 Encode (BRC encoding)
@@ 5.2 Encoding
The encoding procedure of BRC takes a binary string $\mathbf{w}$ as input and outputs a break-resilient codeword, and it is provided in Algorithm 1. At high level, $\mathbf{w}$ will be converted to a sequence of shorter distinct binary strings with the method introduced in the previous section, and their order will be recorded and then protected using a Reed-Solomon code. Synchronization issues among the symbols of the Reed-Solomon code will be resolved using the RLL and MU techniques mentioned earlier.
In detail, let binary string $\mathbf{w}\in{\{0,1\}}^{k}$ be an information word, where
@@
and $l$ and $m$ are positive integers such that $m\geq{\lceil 2\log l\rceil}+2$. The encoding of $\mathbf{w}$ begins by prepending it with $\mathbf{p}^{(0)},\mathbf{p}^{(1)},\cdots,\mathbf{p}^{(t-1)}$, resulting in
@@
where $\mathbf{p}^{(i)}$ is the binary representation of $i\in[0,t-1]$ in $m$ bits (line 3). Then, as shown in line 4, the resulting string $\mathbf{u}$ is fed into the function d-encode (which is implemented in Algorithm 3) and mapped to an array of $l$ pairwise distinct binary strings of length $m$, i.e.,
@@
$\texttt{next}[\mathbf{s}]=\begin{cases}\mathbf{u}_{i+1}~{}&\text{if}~{}\mathbf{s}=\mathbf{u}_{i}~{}\text{for}~{}i\in[0,l-2],\\ \mathbf{s}~{}&\text{otherwise.}\end{cases}$
Note that the value $\texttt{next}(\mathbf{s})$ is well defined, since $\mathbf{u}_{i}\neq\mathbf{u}_{j}$ for every $i\neq j$ by the pairwise distinct property of dStrings. It is also worth noting that the mapping from dStrings to next is injective; one may recover dStrings from next by observing every key $\mathbf{r}$ such that $\texttt{next}(\mathbf{r})\neq\mathbf{r}$, and connecting every two $\mathbf{r}_{a},\mathbf{r}_{b}$ of them if $\texttt{next}(\mathbf{r}_{a})=\mathbf{r}_{b}$.
@@
For every $i\in[t]$$\textsc{rll}(\mathbf{d}_{i})$ does not contain any codeword of $\mathcal{C}_{\text{MU}}$.
@@ 5.3 Decoding
The decoding algorithm may fail to find a non-marker (i.e., $\mathbf{u}_{i}$ for $i\geq t$) in the information region if a break occurs inside it. Hence, the algorithm fails to identify at most $t_{1}$ non-markers. Notice that, failing to capture a non-marker $\mathbf{u}_{a}$ affects at most two key-value pairs of next, i.e., $\texttt{next}[\mathbf{u}_{a-1}]$ and $\texttt{next}[\mathbf{u}_{a}]$, and hence next and approxNext differ in at most $2t_{1}$ positions. ∎
The decoding algorithm proceeds to correct the constructed key-value store approxNext to next, i.e., the key-value store generated in Algorithm 1 from $\mathbf{w}$, using the collected redundancy strings and a standard Reed-Solomon decoder (line 10). Note that the missing redundancy strings are erasures (line 3), since their locations are known. The success of the decoding process is guaranteed in the following theorem.
@@
Alg. 2, line 10 outputs the correct key-value store next.
@@
Meanwhile, Lemma 3 implies that the number of key-value pairs in which next and approxNext differ is bounded by $2t_{1}$. Hence, the systematic part of the constructed codeword has at most $2t_{1}$ errors. Recall that the encoding process generated $4t$ redundancy strings (symbols in $\mathbb{F}_{2}^{m+1}$), and as a result, the decoding in line 10 is guaranteed to be successful since
@@
Next, the array dStrings is obtained from next (line 11), and the function d-decode (which is implemented in Algorithm 4 in the appendix) is employed to recover $\mathbf{u}$. Recall that $\mathbf{u}=\mathbf{p}^{(0)}\circ\cdots\circ\mathbf{p}^{(t-1)}\circ\mathbf{w}$, and hence the decoding procedure concludes by returning the $l\cdot m$ rightmost bits of $\mathbf{u}$ (line 13).
Finally, the decoding procedure implicitly implies that the break-resilient code is also lost-resilient up to a certain threshold, shown as follows.
Corollary 1.
The break-resilient code tolerates the loss of fragments that are shorter than $82m+2{\lceil\log m\rceil}+$ 8 bits.
Proof.
Alg. 2 does not use any fragments that is shorter than twice of the length of a codeword in $\mathcal{C}_{\text{MU}}$, which equals to $82m+2{\lceil\log m\rceil}+$ 8. ∎
@@ 6 Implementation
We implemented a prototype of SIDE and conducted experiments to validate its applicability.
6.1 Secure Information Embedding
In prototyping our design shown in Figure 7, we employ a Creality Ender 3 3D printer (Figure 10(a)), and a Raspberry Pi 3B board (with OP-TEE V3.4 support enabled in its ARMv8-A architecture, and Raspbian Linux 4.14.98-v7 installed for the normal world) to serve as the control board. We adopt a bit embedding method by varying layer thickness, with a slight modification to the methods presented in Section 2.2444 Our proposed framework are not inherently restricted to any particular TEE implementation, printer model, and bit embedding technique, as long as those bits are manifested as physical elements on the printing direction. The choices are made for the sake of experiments. .
As shown in Figure 8, every bit in the codeword corresponds to an interval of $0.36$ millimeters along the printing direction. If value of the bit is 00, then the corresponding interval consists of three layers of $0.12$ millimeters. Otherwise, the corresponding interval consists of two layers of $0.18$ millimeters. These values are chosen to maximize information density, i.e., the number of bits that can be embedded per unit length, with respect to the minimum layer thickness ($0.12$ millimeters) offered by the Ender 3 printer used in the experiments.
The choice of this embedding method is based on the argument that any practical bit embedding method should satisfy the following three properties.
@@ 6.1 Secure Information Embedding
The readability requirement implies economical reading of bits from the broken object with respect to a given use case, which will be illustrated in the next section. The robustness requirement states that the embedded information must be sufficiently robust against adversarial tampering, within a certain bound. For example, the surface of the printed object may be easily abraded and as a result, embedding information on the surface is not robust. We note that layer thickness is an intrinsic property of the printed object, and cannot be easily altered, nor it is affected by breaking or scraping the object.
@@
The obliviousness requirement decouples the embedding method both from the information to be embedded and from the 3D models. Examples which violate the obliviousness requirement include representing $1$’s by thick layers and representing 00’s by thin layers; this will cause codewords with different number of 1’s to have embeddings of different lengths. Moreover, embedding bits by altering the geometry of the printed object is also considered inappropriate, as it may significantly affect its functionality. Note that for any embedding method, imposing some requirement on the dimension of the 3D model is inevitable (e.g., that its size is sufficient for embedding a certain number of bits), and we do not consider this as a violation of the obliviousness requirement. In our embedding method, as the thickness of layers add up to $0.36$ millimeters in embedding either 00 or $1$, the situation of height discrepancy is avoided, and the obliviousness requirement is satisfied. An example of printed objects with an embedded codeword is shown in Figure 9(a).
@@
Figure 8: Demonstration of embedding $01101001$ by varying layer thickness, in which each 00 is represented by three layers of $0.12$ millimeters, and each $1$ is represented by two layers of $0.18$ millimeters. The length required for embedding one bit is irrelevant to its value.
Our development is heavily based on the Klipper open-source project [Klipper] and t43 [t43]. The former is a 3D printer firmware known for offering high precision stepper movements with supports to printers with multiple micro-controllers, and is suitable for running on low-cost devices such as Raspberry Pi. The latter is an open-source slicer program with basic functionalities and is purely written in C, making it suitable for trusted environments with restricted memory space and limited language support.
Our implementation follows the design shown in Figure 7. It involves a host-app that runs in the normal world, and four functional modules (codeclayer-gentoolpath-gen, and firmware) that runs in ARM TrustZone. It is worth noting that the printing process involves looping through every bit of the codeword $\mathbf{c}$. During each repetition, the corresponding layers of the bit are sliced and a G-code file is generated accordingly. This design avoids the potentially overlarge intermediate data (e.g., G-code of the entire model) that may exhaust the limited memory space of ARM TrustZone.
The host-app provides an interface for the printer users. It is developed on top of the Kilpper command line interface, and allows users to configure the connected printer and submit 3D model files for printing. Written in Python, the host bridges the normal world with the trusted world using C Foreign Function Interface (CFFI).
The codec implements Algorithm 1 in C. It takes the the to-be-embedded information $\mathbf{w}$ and outputs a break-resilient codeword $\mathbf{c}$. The codec is a necessary component of the SIDE backend in the case where $\mathbf{w}$ is unknown at the manufacturing of the 3D printer, e.g., geolocation of the printing. Otherwise, the codeword may be simply generated in advance and hardcoded in the printer.
The layer-gen is developed on top of t43 and performs segmentation of the model file. In contrast to a normal slicer program that segments a 3D model altogether, layer-gen is called repeatedly during the examination of bits of the codeword $\mathbf{c}$. During each repetition, the layer-gen first determines the heights of the next two (or three) layers from the value of the current bit of $\mathbf{c}$. For each layer, the layer-gen creates a diagram representing the cross section image of the model. The diagram is generated from the triangles in the model (mesh) that intersect with the corresponding layer height, and is represented by a polygonal chain.
The toolpath-gen is similarly developed on top of t43. During each repetition, the toolpath-gen picks up the diagrams with their corresponding heights from layer-gen, and generates nozzle toolpath (G-code) used to manufacture them.
The firmware performs the parsing of G-code generated from toolpath-gen. It is a collection of core functionalities provided by Klipper, including the computation of precise nozzle movement and the generation of stepper events. The stepper events are then converted to signals passed to the printer hardware. Note that the benign users are oblivious to migration from REE to TEE since it does not change the parsing results.
Figure 9: (a) A cuboid with an embedded BRC codeword. The codeword creates grooves on the outside of the object that are visible to the naked eye. For simplicity, we read the codeword using a microscope which detects light reflected from the top of each layer (Fig. 11). In cases where the outside grooves are not visible (e.g., scraping, filing, or coating), one can use a CT scanner to detect layer thickness. (b) A cuboid broken by hand to $3$ fragments. Indeed, breaks occur naturally along cross-sections of the printing axis.
6.2 Secure Information Extraction
The extraction of information begins by scanning the surface of the fragments of the printed object (Figure 9(b)) with a Leica S9D microscope, which provides a high resolution that captures their rugged surface.